Suppose a perfectly competitive firm with total cost function given as: TC = 400+20Q-2Q²+2/3Q³


1
Expert's answer
2021-10-13T10:16:57-0400

"MC = \frac{dTC}{dQ}"


"MC = 20-4Q+Q^2"


"Average cost = \frac{TC}{q}"

"= \frac {400+20Q-2Q^2+\frac{2}{3}Q^3}{Q}"


"Average cost = \frac{400}{Q}+20-2Q+\frac{2}{3}Q^3""AVC(Q)= 20-4Q+Q^2"


"AFC(Q)= \frac{400}{Q}"